OpenStudy (lifeisadangerousgame):

How do I find how many solutions are there to the following system of equations? 2x – y = 2 –x + 5y = 3

OpenStudy (lifeisadangerousgame):

What I've done so far: 5(2x – y = 2) ----> 10x - 5y = 10 (–x + 5y = 3) ---> -x + 5y = 3 Cross out the 5y's.... 10x = 10 -x = 3 then I add so 9x/9 = 13/9

OpenStudy (ajprincess):

so x=13/9 apply this value in ur first eqn and find the value of y.

OpenStudy (lifeisadangerousgame):

ohhh let me do it out one second please :P

OpenStudy (ajprincess):

oh k. I will wait.

OpenStudy (lifeisadangerousgame):

Ah. idk how to do it :/

OpenStudy (ajprincess):

2x-y=2 2*13/9-y=2 nw can u do?

OpenStudy (lifeisadangerousgame):

OH that's what I did wrong! I made it "y" and not x!

OpenStudy (ajprincess):

sorry. i didn't get u

OpenStudy (lifeisadangerousgame):

No you're good :P I got y = 8.9 is that correct

OpenStudy (ajprincess):

it has to b 0.89

OpenStudy (lifeisadangerousgame):

sorry, moved the decimal :P

OpenStudy (lifeisadangerousgame):

So A. 0 B. 1 C. 2 D. infinitely many

OpenStudy (lifeisadangerousgame):

Those are teh answers I can choose....how can I find which one it is?

OpenStudy (ajprincess):

only 1 solution

OpenStudy (lifeisadangerousgame):

Could you explain to me why?

OpenStudy (ajprincess):

R u familiar with graphs?

OpenStudy (lifeisadangerousgame):

Yes, and I thnik I found it in my book :P Thanks though

OpenStudy (lifeisadangerousgame):

I hate graphs :P

OpenStudy (lifeisadangerousgame):

Ooh I get it! Thanks so much :D

OpenStudy (ajprincess):

sure u got it. if nt I will explain. no problem

OpenStudy (lifeisadangerousgame):

Nope, I think I'm good :) Would you mind helping me with the rest?

OpenStudy (ajprincess):

ha k. sure post them.

OpenStudy (lifeisadangerousgame):

Thanks!

OpenStudy (lifeisadangerousgame):

Write a system of equations to solve the following problem. Let c be the number of child tickets and a be the number of adult tickets. Each child ticket for a ride costs $3, while each adult ticket costs $5. If the ride collected a total of $115, and 33 tickets were sold, how many of each type of ticket were sold? My answer: 3c + 5a = 115 a + c = 33

OpenStudy (ajprincess):

u r right.

OpenStudy (lifeisadangerousgame):

Thanks! I appreciate you helping me :)

OpenStudy (ajprincess):

u r welcome. :)